OpenStudy (anonymous):

calculus question! prove that lim n-> infinity (2^n n!)/n^n = 0.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} 2^n n!\div n^n =0\]

OpenStudy (turingtest):

prove\[\lim_{n\to\infty}{2^nn!\over n^n}=0\]ha, and you have another good problem going as well, why not focus on that one first while I feebly attempt this one

OpenStudy (turingtest):

log tricks?

OpenStudy (zarkon):

show \[\sum_{n=1}^{\infty}(2^n n!)/n^n\] converges

OpenStudy (turingtest):

root test then?

OpenStudy (zarkon):

ratio test works nice

OpenStudy (anonymous):

can you show me the steps? :0

OpenStudy (zarkon):

it is also easy to get an upper bound for (2^n n!)/n^n that also goes to zero

OpenStudy (turingtest):

\[a_n={2^nn!\over n^n}\]\[a_{n+1}={2^{n+1}(n+1)!\over (n+1)^{n+1}}\]\[\large \lim_{n\to\infty}|{a_{n+1}\over a_n}|=\lim_{n\to\infty}\frac{2n^n}{(n+1)^n}=2e^{\lim_{n\to\infty}n\ln{n\over n+1}}\]just looking at the exponent and using log properties we have\[\lim_{n\to\infty}n\ln{n\over n+1}=...?\]@Zarkon what would you do now? n=1/t ?

OpenStudy (zarkon):

\[\frac{n^n}{(n+1)^n}=\left(\frac{n}{n+1}\right)^n\] \[=\left(\frac{n+1-1}{n+1}\right)^n=\left(1-\frac{1}{n+1}\right)^n\] \[=\left(1-\frac{1}{n+1}\right)^{n+1}\left(1-\frac{1}{n+1}\right)^{-1}\] \[\to e^{-1}\cdot 1=e^{-1} \text{ as }n\to\infty\]

OpenStudy (turingtest):

oh that's just so clever it makes me sick ;) Thanks you master Zarkon!

OpenStudy (zarkon):

I guess we should point out to the OP that if you have a series that converges then the sequence converges to zero. Thus we have shown that \(\displaystyle \lim_{n\to\infty}{2^nn!\over n^n}=0\)

OpenStudy (anonymous):

\[(\frac{n}{n+1})^n=(\frac{n+1}{n})^{-n}=(1+\frac{1}{n})^{-n}=e^{-1}\] just thought i would put my two cents in